OpenStudy (firechild17):

Select the equation of the line that passes through the point (5, 7) and is perpendicular to the line x = 4. y = 4 x = 5 y = 7 x = 7

OpenStudy (firechild17):

@OtherWorldly i think this is A

RhondaSommer (rhondasommer):

well, it would be B actually. :)

OpenStudy (firechild17):

oh thanks

OpenStudy (otherworldly):

yes it would be B

RhondaSommer (rhondasommer):

if you have a straight line, where x=4 then all of your coordinates are going to be (4, 1) (4,2) etc. because you need it to be parrallel, then you want it to be the same thing. the coordinate is (5,7) so you want all the coordinates to be (5,1) (5,2) etc. makind x=5 :)

OpenStudy (amorfide):

you have two sets of points (5,7) and x=4  to be perpendicular, it must intersect at a right angle to the line x=4 so the standard point-slope form is y=mx+b we have to work otu b now we have (5,7), so y=7, x=5 7=m(5)+b we know that m is the slope, which is 0, to be perpendicular to x=4 7=0+b b=7 so since m=0 we have y=b so y=7

RhondaSommer (rhondasommer):

OH GOODNESS! I thought it wanted parallel. lel y=7 would be the correct answer. my bad.

RhondaSommer (rhondasommer):

@firechild17

OpenStudy (firechild17):

so it was b? @RhondaSommer

RhondaSommer (rhondasommer):

no. the answer is C. You want the y coordinate (5,7) (7 being the y coordinate. :)